OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3? Answer choices good luck! x-y+ 2 = 0 x+y+ 2 = 0 x+y- 2 = 0

OpenStudy (jack1):

b

OpenStudy (jack1):

no, sorry, c

OpenStudy (jack1):

x + y = 3 is the same as y = -x +3 runs parallel to y = -x +2